The area of ractangle is 8 2/3 sq. cm. If it's length is 5 7/9 cm, find it's perimeter
Answer by CubeyThePenguin(3113)

width = (area)/(length) = 1 1/2 cm

perimeter = 2(1 1/2) + 2(5 7/9) = 14 5/9 cm